# CORE - A Cell-Level Coarse-to-Fine Image Registration Engine for Multi-stain Image Alignment

### v1

@@ 4.4 Discussion
We evaluated the performance and time efficiency of nuclei detection using classical image processing methods in comparison with deep learning-based approaches. Although deep learning models require annotated nuclei data specific to each dataset, most slides used for registration lack such detailed annotations. To ensure generalisability across datasets, we therefore developed a classical machine learning-based nuclei detection pipeline.
Table A.9 summarises the processing time for each step in the CORE method on both the CPU and GPU platforms. Finally, we investigate the effect of regularisation on accurately preserving tissue morphology during deformation. We conducted extensive ablation studies to evaluate the robustness of our method and to understand the influence of key hyperparameters. Our approach outperforms existing methods in terms of the relative target registration error (rTRE), reinforcing the value of combining structural deep features with biological landmarks. This dual-level strategy is particularly beneficial in cases where modality-specific artefacts or variations, such as uneven staining, fixation-induced deformation, or differences in section thickness, can confound intensity-based methods.
However, there are limitations to consider; the fine registration stage depends on accurate nuclei detection, which can be hindered by overlapping cells, staining artefacts, or suboptimal segmentation in dense tissue regions. A key limitation of our approach is that the fine, shape-aware point-set registration stage is more time-consuming than coarse-level techniques. Nevertheless, this stage delivers substantial performance gains, achieving highly precise nuclei-level alignment for re-stained slides and fine tissue-level alignment for consecutive sections. The overall runtime could be further reduced by integrating faster deep learning models for nuclei detection, thereby enhancing computational efficiency. Moreover, while our feature extractor generalises well across modalities, performance may further improve with stain-specific fine-tuning or the integration of self-supervised domain adaptation techniques. Another consideration is the potential computational burden of full-resolution deformable registration, especially for high-resolution WSIs. Although our use of hierarchical processing mitigates this to some extent, further acceleration, possibly via parallel optimisation, could enhance scalability.
@@ 5 Conclusions
In the fine registration stage, nuclei are detected in both fixed and coarsely registered slides. The shape-aware point set registration is performed at full resolution, followed by deformable alignment using the CPD algorithm to achieve cellular-level precision.
@@ Appendix A Supplementary Details
Staining
Tissue samples are typically prepared using a range of staining techniques to highlight different cellular structures and biomarkers. Common methods, such as H&E, mIHC, and mIF, are used to reveal tissue characteristics that are essential to identify abnormalities and disease-related changes.
H&E is the standard stain used, where haematoxylin highlights cell nuclei in blue, and eosin stains cytoplasmic and extracellular components in varying shades of pink, offering a clear view of overall morphology of the tissue.
IHC, on the other hand, uses antibodies conjugated to chromogenic enzymes to target specific proteins, allowing the visualisation of particular cell types or molecular markers, such as immune cells or cancer biomarkers. mIF enables the simultaneous detection of multiple biomarkers through fluorescent labelling, making it especially valuable for profiling the tumour microenvironment or complex cellular interactions.
Fig. A.1 shows comparison of tissue staining techniques: (Left) Hematoxylin staining highlights cell nuclei in blue. (Middle) Immunohistochemistry using an antibody-chromogen system detects a specific protein in brown. (Right) Multiplex immunofluorescence allows simultaneous visualisation of multiple markers, with nuclei staining with DAPI coloured in blue and other markers in green and red.
WSI Registration
Multimodal WSI registration enables more accurate and comprehensive analysis of tissue samples. This alignment allows for the integration of complementary information from different stains, the reconstruction of tissue from serial sections, and the transfer of annotations across slides. Such capabilities are essential for visualising multiple biomarkers simultaneously, the precise comparison and integration of structural and cellular data, identifying affected tissues, detecting tumour markers, and conducting detailed analyses that inform clinical decisions related to diagnosis, prognosis, and treatment planning.
Beyond clinical applications, WSI registration is widely used in analytical and research contexts. It helps in tracking tumour progression and identifying disease-affected regions.
Deep Learning-based Nuclei Detection
For enhanced robustness and accuracy, our architecture integrates a deep learning-based nuclei detection module based on Efficient-MapDe [raza_deconvolving_2019], a neural network specifically designed for cell detection in WSI and further refined in our dataset. Detection is restricted to tissue-containing regions, as determined by the Florence-derived tissue mask. This selective processing avoids background regions, thereby improving computational efficiency.
The detected nuclei are represented as 2D coordinates in the image space, forming point sets that capture the spatial distribution of cellular structures. These serve as landmarks for the subsequent point set registration process. Although this deep learning module is computationally more intensive, averaging 1800 seconds for the largest WSIs on GPU, it significantly improves the accuracy of the registration, particularly in complex regions with histological artefacts or heterogeneous staining.
By incorporating both detection strategies within a unified framework, our system allows users to balance speed and accuracy according to the requirements of their specific application, offering flexibility across a wide range of use cases.
@@
The primary trade-off between classical and deep learning-based approaches lies in computational efficiency. While the classical method is faster, the deep learning model offers improved accuracy at the cost of increased processing time, particularly when applied to WSIs. Both methods produce nuclei centroids as 2D point sets, which serve as the basis for the subsequent point set registration stage.

### v3

@@ 4.4 Discussion
The main reason we developed a machine learning–based nuclei detection approach, rather than a deep learning model, is that deep learning methods require extensive dataset-specific annotated nuclei data which most slides used for registration lack. To ensure generalisability across datasets, we therefore designed a classical machine learning–based nuclei detection pipeline.
Table A.10 summarises the processing time for each step in the CORE method on both the CPU and GPU platforms. Finally, we investigate the effect of regularisation on accurately preserving tissue morphology during deformation. We conducted extensive ablation studies to evaluate the robustness of our method and to understand the influence of key hyperparameters. Our approach outperforms existing methods in terms of the relative target registration error (rTRE), reinforcing the value of combining structural deep features with biological landmarks. This dual-level strategy is particularly beneficial in cases where modality-specific artefacts or variations, such as uneven staining, fixation-induced deformation, or differences in section thickness, can confound coarse-level methods.
However, there are limitations to consider; the fine registration stage depends on accurate nuclei detection, which can be hindered by overlapping cells, staining artefacts, or suboptimal segmentation in dense tissue regions. A key limitation of our approach is that the fine, shape-aware point-set registration stage is more time-consuming than coarse-level techniques. Nevertheless, this stage delivers substantial performance gains, achieving highly precise nuclei-level alignment for re-stained slides and fine tissue-level alignment for consecutive sections. The overall runtime could be further reduced by integrating faster deep learning models for nuclei detection, thereby enhancing computational efficiency, however it will require explicit training/finetuning for each dataset and we have not covered this aspect as part of the study. Moreover, while our feature extractor generalises well across modalities, performance may further improve with stain-specific fine-tuning or the integration of self-supervised domain adaptation techniques. Another consideration is the potential computational burden of full-resolution deformable registration, especially for high-resolution WSIs. Although our use of hierarchical processing mitigates this to some extent, further acceleration, possibly via parallel optimisation, could enhance scalability.
@@ 5 Conclusions
In the fine registration stage, nuclei are detected in both target and coarsely registered source slides. The shape-aware point set registration is performed at full resolution, followed by deformable alignment using the CPD algorithm to achieve cellular-level precision.
@@ Appendix A Supplementary Details
Fig. A.1 shows comparison of tissue staining techniques: (Left) Haematoxylin staining highlights cell nuclei in blue. (Middle) Immunohistochemistry using an antibody-chromogen system detects a specific protein in brown. (Right) Multiplex immunofluorescence allows simultaneous visualisation of multiple markers, with nuclei staining with DAPI coloured in blue and other markers in green and red.
Using the demo:
To facilitate real-time deformation of the source image and simultaneous visualisation of the transformed and target WSIs, we provide a dedicated viewer, TIAViz. The tool is available as part of the TIATOOLBOX library [pocock_tiatoolbox_2022], which can be installed via pip. The demo can be used by following the steps below:
1.
Install TIAViz. Install the TIATOOLBOX package using pip.
2.
Prepare the directory structure. Within the working directory, create two subdirectories named slides and overlays.
3.
Place the required files. Store the source WSI in the slides directory. Store the target WSI together with deformation field in the overlays directory.
4.
Ensure filename consistency. All corresponding files must share the same stem name so that TIAViz can automatically match the source, target, and deformation field.
5.
Launch the viewer. Start the TIAViz interface using the tiavisualize command, either from the command line or within a Python notebook, specifying the paths to the slides and overlays directories.
6.
Registration visualisation. From the slides dropdown in the top-left panel, first select the source slide. After this, choose the deformation field from the overlay dropdown; the deformation will be applied to the source slide within a few seconds. Next, select the target slide from the Overlay dropdown to display it on top of the registered source. Use the right-hand panel to adjust the alpha values of both layers to achieve optimal visualisation of the alignment.
Stain Normalisation:
Stain Normalisation is often required for H&E slides because their colour appearance varies widely across laboratories, scanners, and staining batches, which can affect how prompt-based models such as Florence-SAM [xiao_florence-2_2023] detect tissue boundaries. Normalising H&E reduces this variability and provides a more consistent input for tissue mask extraction. In contrast, IHC slides uses stable chromogens such as DAB that have less colour variability, and convey biologically meaningful intensity differences that normalisation could distort, so stain normalisation is generally unnecessary for IHC.
@@
| AMTRE [m] | AATRE [m] | Avg Time (s) | AMTRE [m] | Avg Time (s) | |||
| CORE | 65.26 | 109.83 | 18 | 6.81 | 18 | ||
| Fine | CORE | 39.20 | 60.21 | 2340 | 2.26 | 2340 | |